The measure of an angle is 50 more than that of its complement. Find the measure of each angle. 70, 20
Answer by Alan3354(69443)

The measure of an angle is 50 more than that of its complement. Find the measure of each angle.
----------
A + C = 90 degs
(C+50) + C = 90
2C + 50 = 90
2C = 40
C = 20
A = 70